A portable ladder is needed to climb onto a rooftop 14 feet high? How long should the ladder be? How far should it's base set away from the building wall?
Answer by oberobic(2304)

Your book must have given you more information on what the proper angle should be for ladder climbing to be safe. Also, how far should the ladder project above the roof so that you are secure and able to hold it as you step from the ladder onto the roof.
.
Assuming you can climb a vertical ladder, then it would need to be 14 feet to climb a 14-ft wall.
Assuming you are fine climbing a ladder that is set 7 feet from the wall (1/2 the height), then you use the Pythagorean theorem.
a = height of the wall
a = 14
b = base of the triangle
b = 7
Pythagoras taught:
c = sqrt(14^2 + 7^2)
c = sqrt(196 + 49)
c = sqrt(245)
c = 15.652
So the ladder needs to be 15.652 feet.
.
That's the decimal answer. However, ladders come in standard lengths in whole feet. If you want to be safe, you need a 20 ft ladder. 16 ft would reach, but it not extend high enough above the roof for you to have a safe handhold when you get off or get on the ladder.
.
Done.